# Hyperelliptic Shimura curves and $L$-functions of central vanishing order at least 3

### v1

@@ 7. An intriguing hyperelliptic Shimura curve of Dembélé
Let $F={\mathbb{Q}}\left(\cos\frac{1}{16}\pi\right)$. In this section, we describe a hyperelliptic Shimura curve $X_{U}$ over $F$ found by Dembélé [4, 5.6]. We first clarify that in [4, 5.6], a Shimura curve is of the form $X_{V}$ with $V$ open compact (instead of compact-modulo-center as in our paper) in $B^{\times}({\mathbb{A}}_{\mathrm{f}})$. And the hyperelliptic curve constructed in loc. cit. is an Atkin–Lehner quotient of such a Shimura curve, which is still a Shimura curve in our sense.
@@
Since the narrow class number of $F$ is 1, $X_{V}$ is geometrically connected and $X_{V}\to X_{V{\mathbb{A}}_{\mathrm{f}}^{\times}}$ is an isomorphism. Here ${\mathbb{A}}_{\mathrm{f}}^{\times}$ is the center of $D^{\times}({\mathbb{A}}_{\mathrm{f}})$. This is the Shimura curve $X_{0}^{D}(1)$ in [4, Introduction, 4.1]. The generator of $D_{\mathfrak{p}}^{\times}/F_{\mathfrak{p}}^{\times}O_{{\mathfrak{p}}}^{\times}\simeq{\mathbb{Z}}/2{\mathbb{Z}}$ normalizes $V_{\mathfrak{p}}F_{\mathfrak{p}}^{\times}$ and thus defines an automorphism of $X_{V{\mathbb{A}}_{\mathrm{f}}^{\times}}$ of order 2. (For example we can take the generator to be the uniformizer of a ramified quadratic extension of $F_{\mathfrak{p}}$ in $D_{\mathfrak{p}}$.) This is the Atkin–Lehner involution $w_{D}$ in [4, Introduction, 5.4]. In particular, the Atkin–Lehner quotient $X_{0}^{D}(1)/\langle{w_{D}}\rangle$ in [4, Introduction, 5.4] is $X_{U}$ with
@@
By [4, 5.3], there are 40 holomorphic Hilbert newforms of level $\Gamma_{0}({\mathfrak{p}})$ and parallel weight 2, and they form 5 Galois orbits of sizes 4, 4, 4, 4 and 24 respectively. Moreover, let $w$ be the Atkin–Lehner involution acting on the $\Gamma_{0}({\mathfrak{p}})$-Hilbert modular forms. Then the newforms in the four Galois orbits of sizes 4 are exactly the ones with $w$-eigenvalue 1-1. Their Jacquet–Langlands correspondences to $D^{\times}$ exactly the ones with $w_{D}$-eigenvalue $1$. So these Jacquet–Langlands correspondences descend to $X_{U,{\mathbb{C}}}$ and span the space of holomorphic differential 1-forms on $X_{U,{\mathbb{C}}}$, which has genus 16. Then by 5.6, we have the following.
@@
Consider the 16 holomorphic Hilbert newforms of level $\Gamma_{0}({\mathfrak{p}})$ and parallel weight 2 with $w$-eigenvalue 1-1. Let $f_{1},f_{2},f_{3}$ be any three of them. Then
@@
Remark 7.3.
(1) Computationally, it is not easy to verify whether the 3rd central derivatives of these $L$-functions vanish due to their large conductors.
(2) 6.2 suggests that we also should check whether the central derivatives of the $L$-functions vanish for single newforms in the theorem. This is easy for people with access to enough computational power.

### v2

@@ 8. An intriguing hyperelliptic Shimura curve of Dembélé
Let $F={\mathbb{Q}}(\zeta_{32}+\zeta_{32}^{-1})$, the maximal totally real subfield of the cyclotomic field of 32nd roots of unity. In this section, we describe a hyperelliptic Shimura curve $X_{U}$ over $F$ found by Dembélé [3, 5.6]. We first clarify that in [3, 5.6], a Shimura curve is of the form $X_{V}$ with $V$ open compact (instead of compact-modulo-center as in our paper) in $B^{\times}({\mathbb{A}}_{\mathrm{f}})$. And the hyperelliptic curve constructed in loc. cit. is an Atkin–Lehner quotient of such a Shimura curve, which is still a Shimura curve in our sense.
@@
Since the narrow class number of $F$ is 1, $X_{V}$ is geometrically connected and $X_{V}\to X_{V{\mathbb{A}}_{\mathrm{f}}^{\times}}$ is an isomorphism. Here ${\mathbb{A}}_{\mathrm{f}}^{\times}$ is the center of $D^{\times}({\mathbb{A}}_{\mathrm{f}})$. This is the Shimura curve $X_{0}^{D}(1)$ in [3, Introduction, 4.1]. The generator of $D_{\mathfrak{p}}^{\times}/F_{\mathfrak{p}}^{\times}O_{{\mathfrak{p}}}^{\times}\simeq{\mathbb{Z}}/2{\mathbb{Z}}$ normalizes $V_{\mathfrak{p}}F_{\mathfrak{p}}^{\times}$ and thus defines an automorphism of $X_{V{\mathbb{A}}_{\mathrm{f}}^{\times}}$ of order 2. (For example we can take the generator to be the uniformizer of a ramified quadratic extension of $F_{\mathfrak{p}}$ in $D_{\mathfrak{p}}$.) This is the Atkin–Lehner involution $w_{D}$ in [3, Introduction, 5.4]. In particular, the Atkin–Lehner quotient $X_{0}^{D}(1)/\langle{w_{D}}\rangle$ in [3, Introduction, 5.4] is $X_{U}$ with
@@
By [3, 5.3], there are 40 holomorphic Hilbert newforms of level $\Gamma_{0}({\mathfrak{p}})$ and parallel weight 2, and they form 5 Galois orbits of sizes 4, 4, 4, 4 and 24 respectively. Moreover, let $w$ be the Atkin–Lehner involution acting on the $\Gamma_{0}({\mathfrak{p}})$-Hilbert modular forms. Then the newforms in the four Galois orbits of sizes 4 are exactly the ones with $w$-eigenvalue 1-1. Their Jacquet–Langlands correspondences to $D^{\times}$ exactly the ones with $w_{D}$-eigenvalue $1$. So these Jacquet–Langlands correspondences descend to $X_{U,{\mathbb{C}}}$ and span the space of holomorphic differential 1-forms on $X_{U,{\mathbb{C}}}$, which has genus 16. Then by 5.6, we have the following.
@@
Consider the 16 holomorphic Hilbert newforms of level $\Gamma_{0}({\mathfrak{p}})$ and parallel weight 2 with $w$-eigenvalue 1-1. Let $f_{1},f_{2},f_{3}$ be any three distinct ones among them. Then
@@
1.1 follows from 8.2.